OpenStudy (anonymous):

Natalie can deliver her newspapers in 83 minuites. It takes Jake 65 minuites to do the same route. How long would it take them to deliver the newspapers if they work together?

hero (hero):

\[\frac{N \times J}{N+J} = t\]

OpenStudy (anonymous):

Seriously? it's that easy?

hero (hero):

Yup

OpenStudy (anonymous):

I got for the top 5395 and for the bottom 146

OpenStudy (anonymous):

It would take them 36.95 minutes?

hero (hero):

Yes

OpenStudy (anonymous):

And if I round it to one decimal place it would be 37 Minutes. correct?

hero (hero):

approximately, yes

OpenStudy (anonymous):

Cool thanks Hero.

OpenStudy (anonymous):

That helps. Thanks a lot.

hero (hero):

\[\frac{abc}{ab +bc + ac}\]

hero (hero):

sorry, it's the other way around.

hero (hero):

Just remember that multiplication goes at the top and addition goes at the bottom.

OpenStudy (anonymous):

ok